OpenStudy (anonymous):

If L is the line through the point (- 3, 5) having a slope of , complete the equation of L. y = (2/3)x + 7 (2/3)x + (47/9) (2/3)x - (1/3)

OpenStudy (darkprince14):

what is its slope?

OpenStudy (anonymous):

y-y1 = m(x-x1) where (x1,y1) = (-3,5) and m= 2/3 Answer is A